June 20, 2013 Solenoid Meeting

From GlueXWiki
Jump to: navigation, search

Thursday June 6,2013 at 11am At Jefferson Lab, the meeting will be held in room CC F326


C. Rode, T. Whitlatch, E. Wolin, B. Flora (via phone for short duration..sorry), E. Chudakov, J. Creel, G. Young, P. Ghoshal


  1. Complete 140A procedure (Rev C) - George (this is in review now with some small changes needed)Media:140A before changes for PLC.pdf
  2. Review 140A procedure and comment - all
  3. Update hardware QD with dial pot and PLC with inductance algorithm - Elliot Media:Proposed PLC QD Changes.pdf
  4. Clean up refrigerator and get ready for cooldown - Jonathan
  5. Determine if turn to turn short could have caused quench in coil 1A - Eugene
  6. Write up quench analysis, dump resistor and max current rationale for review - George Media:Narrative.pdf
  7. Review previous low current ramp for voltage spikes and perform additional low rate ramping if necessary - Giles
  8. Write up high current portion of procedure (Rev D) - George
  9. Balance Power supply for 25H (if necessary based on item 7)- Giles
  10. Implement fast dump wire connections from refrigerator - Jonathan/Elliot - complete
  11. Calculate what temperature is needed to run at 1500A and 1360A with 0.9K margin. Media:Quench Margins.pdf
  • What Current do we map the magnet at?
  • Is it worth spending all this effort on the QD if it will be changing for production?

Useful Info

  • Documents available at M:\halld_engineering\Solenoid\Commissioning\Response to Quench

Meeting Minutes

Briefly went over last week's meeting minutes and path forward. There has been alot of email chatter on the quench detector configuration. Bob Flora is happy with the current plan on using the software QD as the primary and the hardware as the secondary possible at a higher trip value. The PXI changes supporting this are done and the PLC changes should be complete early next week. George will reissue the 140A procedure by tomorrow morning. High current preliminary procedure and the updated quench write up by COB Monday. We will change out the dial pots on the hardware QD with a 10 turn balance pot and a 3 3/4 turn gain pot early next week. Jonathan's crew is still working on the refrigerator and anticipates he will be ready for cooldown by next Wednesday. Need to make a decision by Tuesday as whether or not to start cooldown of the refrigerator on Wednesday since Patrizia will be away from the lab starting Wednesday. Discussion of turn to turn shorts determined that it is not worth the effort to do an R&D run to determine what can be detected. It was suggested that we use a lower ramp rate at high currents to reduce any effects and George will incorporate this into the high current run procedure. George and crew will continue with ramp rate test at low current to look at the effects on voltage spikes. It was a general consensus that if the voltage spikes are small enough, the 25 Henry tuning hardware will not be installed until after mapping. The current thinking for high current is to go to 1360A with dewar installed (from some lower current)and then go to 1350 and map the beamline location only. Then reduce to 1300A and fully map the magnet. When the dewar runs out it will not cause any pressure spikes. We will assume the pressure spike from connecting the dewar will bring the temp to 5K (very conservative). Probir will determine the max current we can run at 5K with a .9K margin per Alcorn's spec. It was also sgreed upon that it would be prudent to go to a slow dump if the PXI heartbeat dies.

New Action Items

  1. Release 140A procedure (Rev C) - George
  2. Update hardware QD with new 10 turn pot and PLC with inductance algorithm - Elliot
  3. Clean up refrigerator and get ready for cooldown - Jonathan
  4. Finalize write up of quench analysis, dump resistor and max current rationale - George
  5. Continue low current ramp tests to categorize voltage spikes - Giles
  6. Write up high current portion of procedure (Rev D) - George
  7. Balance Power supply for 25H (if necessary based on 140A tests)- Giles
  8. Calculate what current we can safely run at at 5K 0.9K margin.
  9. Set up meeting with Patrizia Tuesday for cooldown decision.

Path Forward

  1. Complete the 140A procedure including the quench balance and turn to turn short check not completed in the 15A test (to be completed and signed off by June 21)
  2. Continue repair of refrigerator (to be completed by June 25)
  3. Update preliminary analysis of quench and rationale for dump resistor and max current (June 25)
  4. Initial review of updated write up by Bill Schneider and Bob Flora (June 26)
  5. Start refridge cooldown (to be started June 27th if agreed upon)
  6. Complete and sign off maximum current plan and quench analysis(to be completed by June 29)
  7. Start cooling magnet - July 8
  8. Perform 140A testing ( to be completed by July 21)
  9. Perform High current commissioning ( to be completed by July 23)
  10. Map magnet (to be completed by July 28)